OpenStudy (anonymous):

f(x) varies inversely with x and f(x) = –15 when x = 30. What is the inverse variation equation? a) f(x)= -5/x b) f(x)= -2/x c) f(x)= -90/x d) f(x)= -450/x

OpenStudy (anonymous):

the equations are already in inverse form... if you switch x and y... x = a/y. Solve for y xy = a y = a/x The form doesn't change!! So you can just plug x and y into those equations to see which one works.

OpenStudy (anonymous):

so its d i believe? @piglet9

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thanks!